The length of a rectangle is 7 in. more than twice its width. If the perimeter of the rectangle is 32 in., find the width of the rectangle.
Answer by tutorcecilia(2144)

Use the formula for the perimeter of a rectangle:
P = 2(length+width)
P=32
Length=width + 7
Width = w
.
Plug-in the values and solve:
P=2(l+w)
32=2(w+7+w)
32=2(2w+7)
32/2=2(2w+7)/2
16=2w+7
16-7=2w
9=2w
9/2=2w/2
4.5=w
.
Check by plugging all of the values back into the original equation:
32=2(4.5+7+4.5)
32=32 [checks out]